OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that satisfies the following condition. slope 9 and passes through (2, 22) solve!

OpenStudy (texaschic101):

y = mx + b slope(m) = 9 (2,22)...x = 2 and y = 22 now we sub and solve for b, the y intercept y = mx + b 22 = 9(2) + b 22 = 18 + b 22 - 18 = b 4 = b so your equation is : y = 9x + 4 any questions ?

OpenStudy (anonymous):

the answers in this question are ; a.y = 9x + 4 b.y = 9x - 22 c.y = 22x - 2 d.y = 22x + 4 just so you know :)

OpenStudy (anonymous):

No thank you so much dude you don't know how much I was needing that !! :D :D

OpenStudy (texaschic101):

no problem :) Happy to help